Grade 12th passOrganic Chemistry

2g of mixture of CO and CO2 on reaction with excess I2O5 produced 2.54g I2.Whay would be the mass % of CO2 in the original mixture?

Somasree Dey
8 Years agoGrade 12th pass

1 Answer

Arun
8 Years ago
Dear student
 
5 CO + I2O5 ------> 5 CO2 + I2
 
1 mol of I2 = 1 mol of I2O5 = 5 mol of CO
 
Hence mol of CO = 5 * 2.54/254 = 0.05
 
Mass of CO = 0.05 * 28 = 1.4 gm
 
Mass of CO2 = 2 - 1.4 = 0.6 gm
 
Mass % of CO2 = 30%
 
 
Regards
Arun (askIITians forum expert)